OpenStudy (anonymous):

chose the point slope form of the equation below that represents the line that passes through the points (-6,4) (2,0)

Parth (parthkohli):

\[y-y_1 = m(x - x_1)\]This is the point-slope form. Okay?

OpenStudy (anonymous):

after i do that then what??

Parth (parthkohli):

Then you substitute the information.

OpenStudy (anonymous):

what i did does not match any of the options, do you mind helping me out more? sorry im really bad at math

Parth (parthkohli):

Can you find the slope of these two points?

OpenStudy (anonymous):

what do you mean?

Parth (parthkohli):

You don't know what the slope is? if there are two points \((x_1,y_1)(x_2,y_2)\), then the slope is:\[ {y_2 - y_1 \over x_2 - x_1}\]

OpenStudy (anonymous):

i got -4/8 or -2/4?

Parth (parthkohli):

That's correct. You now have:\[ y - y_1 = {-2 \over 4}\left(x - x_1 \right)\]Remember that the slope is denoted by \(m\).

Parth (parthkohli):

Can you tell me what \(y_1\) and \(x_1\) are?

OpenStudy (anonymous):

umm -6 and 4?

Parth (parthkohli):

Yep.\[x_1 = -6\\ y_1 = 4 \]Substitute that info in the equation:\[y-y_1 = {-2 \over 4}\left(x - x_1 \right) \implies y - 4 = {-2 \over 4}(x - (-6))\]

Parth (parthkohli):

Believe me - you could simplify that much more.

OpenStudy (anonymous):

i know i have to distribute the fraction but how do i do that?

Parth (parthkohli):

Try the first step, which is simplifying \(x-(-6)\).

OpenStudy (anonymous):

i got the awnser! i submit my assignment and i got it right! thank you sooo much for the help!!!